OpenStudy (anonymous):

An advertising agency conducted a survey and found that the number of units sold, N, is related to the amount a spent on advertising (in dollars) by the following formula: N = 4,500 + 200 ln a (a ≥ 1) How many units are sold after spending $1,000? (Round your answer to the nearest whole number.)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

N = 4500 + 200*ln(a) N = 4500 + 200*ln(1000) ... plug in a = 1000 N = ???

OpenStudy (anonymous):

and i did that and got 5882? that was the procedure i performed just wanted to check answer

OpenStudy (anonymous):

thanks for your help again!

jimthompson5910 (jim_thompson5910):

you got the right answer